# Online Multivariate Changepoint Detection: Leveraging Links With Computational Geometry

### v2

@@ H.3 Some peeling bounds
Assume $\varepsilon_{t}$ are i.i.d. Gaussian with identity variance matrix. For $\alpha$ in $(0,1)$ we have:
  • Sparse statistic
    $\mathbb{P}(\exists s,n\in\mathbb{N}\ |\ S_{s,n}>8\log(np/\alpha))\leq\alpha.$ (24)
  • Dense statistic Taking $x_{n,\alpha}=4\log(n)-log(\alpha)+1,$ we have
    $\mathbb{P}(\exists s,n\in\mathbb{N}\ |\ D_{s,n}>p+2\sqrt{px_{n,\alpha}}+2x_{n,\alpha}))\leq\alpha\,;$ (25)
    $\mathbb{P}(\exists s,n\in\mathbb{N}\ |\ D_{s,n}<p-2\sqrt{px_{n,\alpha}}))\leq\alpha.$ (26)
@@
For the dense case, the proof follows the peeling argument of Lemma 8 in [65] replacing the tail bound from their sub-Gaussianity assumption with the upper and lower bound on the tail probability of a chi-square distribution with $p$ degrees of freedom from Lemma 1 of [66]. For any non-decreasing sequence $b_{n}$ applying a union bound (that is for any $n$ events $A_{i}$ with $i$ in $\{1,\ldots n\}$ $\mathbb{P}(\cup_{i}A_{i})\leq max_{i}n\mathbb{P}(A_{i})$) several times we get:
$\mathbb{P}\left(\exists s,n\in\mathbb{N}\ |\ D_{s,n}>b_{n}^{2}\right)$$\leq$ $\sum_{j=0}^{\infty}\mathbb{P}\left(\max_{2^{j}\leq n<2^{j+1}}\max_{1\leq s<n}D_{s,n}>b_{n}^{2}\right)$
$\leq$ $\sum_{j=0}^{\infty}2^{j}\max_{2^{j}\leq n<2^{j+1}}\mathbb{P}\left(\max_{1\leq s<n}D_{n,s}>b_{n}^{2}\right)$
$\leq$ $\sum_{j=0}^{\infty}2^{j}\max_{2^{j}\leq n<2^{j+1}}n\mathbb{P}\left(W^{2}>b_{n}^{2}\right)$
$\leq$ $\sum_{j=0}^{\infty}2^{2j+1}\mathbb{P}\left(W^{2}>b_{2^{j}}^{2}\right),$
where $W^{2}$ has a chi-square distribution with $p$ degrees of freedom. We then take $b_{n}^{2}=(p+2\sqrt{px^{\prime}_{\alpha,n}}+2x^{\prime}_{\alpha,n})$ with $x^{\prime}_{\alpha,n}=2\log(n)+\log(\log(2n))+\log(\log(4n))-\log(2)-2\log(\log(2))-\log(\alpha)$. Using the tail-bound from [66] we get that
$2^{2j+1}\mathbb{P}(W^{2}>b_{2^{j}}^{2})\leq 2^{2j+1}\exp(-x^{\prime}_{\alpha,2^{2j}})$
$\leq$ $\exp((2j+1)\log(2)-2j\log(2)-\log(2)-\log((j+1))-\log((j+2))+\log(\alpha))$
$\leq$ $\frac{\alpha}{(j+1)(j+2)}.$
Summing over all $j$ we get $\mathbb{P}(\exists s,n\in\mathbb{N}\ |\ S_{s,n}>b_{n}^{2})\leq\alpha.$ Out of simplicity we consider $x_{\alpha,n}=4\log(n)-\log(\alpha)+1$ which is larger than $x^{\prime}_{\alpha,n},$ because $2\log(n)\geq\log(\log(2n))+\log(\log(4n))$ for $n\geq 1$, and $1\geq-\log(2)-2\log(\log(2)).$
We get the lower bound (26) similarly considering the lower bound in Lemma 1 of [66].
For the sparse case we apply the peeling argument of Lemma 8 of [65] to each variate and use a union bound overall $p$ variates. ∎
Remark 10.
The proof would still work if we consider a fixed and finite maximum run length, $n_{max}$ say, as for any $j_{max}$ we have $\sum_{j=0}^{j_{max}}\frac{1}{j+1}-\frac{1}{j+2}\leq 1$. Furthermore, as the threshold increases with $j$, with a finite maximum run length the proof would still hold with the fixed threshold $x_{n_{max},\alpha}$.
We now recall a bound of [68] to control the variations of $O_{s,n}^{i}$ (or more generally $\frac{1}{\sqrt{(n-s)}}\sum_{t=s+1}^{n}\varepsilon^{\prime}_{t}$ for $\varepsilon^{\prime}_{t}$ i.i.d Gaussian with variance 1 and mean 0) which is useful to study the detecting delay (see subsubsection H.2.1).
@@
For a given $s$ in $\mathbb{N}$ and $i\in\{1,,\ldots,p\}$,
$\mathbb{P}\biggl{(}\exists n\in\mathbb{N}\ |\ O_{s,n}^{i}>2\sqrt{2\log(\log(3(n-s)))-\log(\alpha)+1}\biggl{)}\leq\alpha.$
@@
We follow the proof of Lemma 5 in [68]. We highlight below the key steps. First, based on their Lemma 4 we have
$\mathbb{P}\biggl{(}\max_{s+2^{j}\leq n<s+2^{j+1}}O_{s,n}^{i}>2x\biggl{)}\leq\exp{(-x^{2}/4)}\,.$ (27)
Second, we apply a union bound over all $j$ :
$\mathbb{P}\left(\exists j\ \bigg{|}\ \sup_{s+2^{j}\leq n<s+2^{j+1}}O_{s,n}^{i}>2\sqrt{\log[(j+1)(j+2)]-\log(\alpha)}\right)\leq\sum_{j=0}^{\infty}\frac{\alpha}{(j+1)(j+2)}=\alpha.$ (28)
In the last step we use $\frac{1}{(j+1)(j+2)}=\frac{1}{j+1}-\frac{1}{j+2}$ and the telescoping nature of the resulting sum. Finally, we note that for any $n$ such that $s+2^{j}\leq n<s+2^{j+1}$ then
$\log((j+1)(j+2))\leq 2\log(\log(3(n-s)))+1.$
Remark 11.
As for Lemma 4 the proof would still work if we consider a fixed and finite maximum run length, $n_{max}$ say, as for any $j_{max}$ we have $\sum_{j=0}^{j_{max}}\frac{1}{j+1}-\frac{1}{j+2}\leq 1$. Furthermore, as the threshold increases with $j$, with a finite maximum run length the proof would still hold considering a fixed threshold $2\sqrt{2\log(\log(3n_{max}))-\log(\alpha)+1}$.

### v3

@@ I.5 Some concentration bounds
Assume $\varepsilon_{t}$ are i.i.d. Gaussian with identity variance matrix. For $\alpha$ in $(0,1)$, taking $x_{n}(\alpha)=4\log(n)-\log(\alpha),$
  • Rank-1 statistic
    $\mathbb{P}\left(\ \exists\ \tau<n\in\mathbb{N}\quad\text{s.t.}\quad S^{\#1}_{\tau,n}>2x_{n}(\alpha)+2\log(2p)\ \right)\leq\alpha.$
  • Rank-p statistic
    $\mathbb{P}\left(\ \exists\ \tau<n\in\mathbb{N}\quad\text{s.t.}\quad S^{\#p}_{\tau,n}>2x_{n}(\alpha)+2\sqrt{px_{n}(\alpha)}+p\ \right)\leq\alpha\,;$
  • Thresholded statistic
    $\mathbb{P}\left(\ \exists\ \tau<n\in\mathbb{N}\quad\text{s.t.}\quad S^{(a)}_{\tau,n}>4x_{n}(\alpha)+6pe^{-a^{2}/8}\right)\leq\alpha.$
  • Rank-s statistic
    $\mathbb{P}\left(\exists\ \tau<n\in\mathbb{N}\ \text{s.t.}\ S^{\#s}_{\tau,n}>2(x_{n}(\alpha)+\log{p\choose s})+2\sqrt{s(x_{n}(\alpha)+\log{p\choose s})}+s\right)\leq\alpha\,;$
@@
For all statistics, the proof follows the peeling argument of Lemma 8 in Yu et al. [2023]. In detail, for some generic statistic $\tilde{S}_{\tau,n}$ and some non-decreasing $b_{n}$, we can bound the probability of the event $\exists\ \tau<n\in\mathbb{N}\ \text{such that}\ \ \tilde{S}_{\tau,n}>b_{n}$ using two union bounds:
$\mathbb{P}\left(\ \exists\ \tau<n\in\mathbb{N}\ \text{such that}\ \tilde{S}_{\tau,n}>b_{n}\right)$$\leq$ $\sum_{j=1}^{\infty}\mathbb{P}\left(\ \exists\ \tau<n\in[2^{j},2^{j+1})\ \text{such that}\ \tilde{S}_{\tau,n}>b_{n}\right)$
$\leq$ $3\sum_{j=1}^{\infty}2^{2j-1}\max_{\begin{array}[]{c@{}c@{}c@{}c@{}c@{}}2^{j}&\leq&n&<&2^{j+1}\\ 1&\leq&\tau&<&n\end{array}}\mathbb{P}\left(\tilde{S}_{\tau,n}>b_{n}\right)$
$\leq$ $3\sum_{j=1}^{\infty}2^{2j-1}\max_{1\leq\tau<2^{j}}\mathbb{P}(\tilde{S}_{\tau,2^{j}}>b_{2^{j}}).$
We get the second inequality bounding the number of couples $(\tau,n)$ such that $\tau<n$ with $n$ in $[2^{j},2^{j+1})$, there are:
$\frac{(2^{j+1}-2^{j})(2^{j+1}-2^{j}-1)}{2}+(2^{j+1}-2^{j})^{2}=2^{j-1}(3\times 2^{j}-1)\,.$
We obtain the last equality using the fact that $n\to b_{n}$ is an increasing bound and that all the $\tilde{S}_{\tau,n}$ have the same distribution, meaning that $n\to\mathbb{P}(\tilde{S}_{\tau,n}>b_{n})$ is non-increasing. To continue, we consider some appropriate concentration bound on all $\tilde{S}_{\tau,n}$ of the form $\mathbb{P}\left(\tilde{S}_{\tau,n}\geq c(x)\right)\ \leq\ e^{-x}$. To be specific:
  • For the rank-1 statistic we take
    $P\left(S^{\#1}_{\tau,n}\geq c(x)\right)\ \leq\ e^{-x},$ (23)
    with $c(x)=2x+2\log(p)+2\log(2)$ which we obtain using a union bound over $p$ univariate Gaussian random variable.
  • For the rank-p and rank-s statistics we take
    $P\left(S^{\#s}_{\tau,n}\geq c(x)\right)\ \leq\ e^{-x},$ (24)
    with $c(x)=s+2\sqrt{s(x+\log{p\choose s})}+2(x+\log{p\choose s})$ which we get from Lemma 1 of Laurent and Massart [2000] and a union bound over all (${p\choose s}$ choices of $s$ dimensions out of $p$.
  • For the threshold statistic we take
    $P\left(S^{(a)}_{\tau,n}\geq c(x)\right)\ \leq\ e^{-x},$ (25)
    with $c(x)=6pe^{-a^{2}/8}+4x$ which we get from Lemma 17 of Chen et al. [2022].
Setting $b_{2^{j}}=c(\tilde{x}_{2^{j}}(\alpha))$, we remove the probability term using these concentration bounds to write:
$\mathbb{P}\left(\ \exists\ \tau<n\in\mathbb{N}\ \text{s.t.}\ \tilde{S}_{\tau,n}>b_{n}\right)\leq 3\sum_{j=1}^{\infty}2^{2j-1}\exp(-\tilde{x}_{2^{j}}(\alpha))\,.$
Considering $\exp(\tilde{x}_{2^{j}}(\alpha))=\frac{3}{\alpha}2^{2j-1}j(j+1)$ we get:
$3\sum_{j=1}^{\infty}2^{2j-1}\exp(-\tilde{x}_{2^{j}}(\alpha))=\sum_{j=1}^{\infty}\frac{\alpha}{j(j+1)}=\alpha\,.$
and we have the value for evaluating $c$ given by:
$\tilde{x}_{n}(\alpha)=2\log(n)-\log(\alpha)+\log(\log_{2}(n)(\log_{2}(n)+1))+\log(\frac{3}{2})\leq 4\log(n)-\log(\alpha)=x_{n}(\alpha)\,,$
using relation $2^{2j}\geq\frac{3}{2}j(j+1)$ on integers for the inequality. As the function $c$ is increasing, we can use the simpler value $x_{n}(\alpha)$ which leads to the proposed result.
We now provide some bounds to control the detection delay, assuming we know the pre-change mean and the threshold is fixed or concave increasing.
@@
Consider a sequence with a change $\delta$ at $\tau^{*}$, that is $y_{t}=\varepsilon_{t}$ for $t\leq\tau^{*}$ and $y_{t}=\delta+\varepsilon_{t}$ for $t>\tau^{*}$ and $\lVert\delta\rVert>0$.
  • Suppose the threshold $c_{n}=c$ is fixed and positive. Then for any
    $n-\tau^{*}\geq\frac{2c-8\log(\alpha/2)}{\lVert\delta\rVert^{2}}=\Delta(\delta,c,\alpha)$
    we have
    $\mathbb{P}\Big{(}\lVert\frac{\sum_{t=\tau^{*}+1}^{n}y_{t}}{\sqrt{n-\tau^{*}}}\|^{2}<c\Big{)}\leq\alpha.$
  • Suppose the threshold $c_{n}=c$ is fixed, positive and larger than $p$. Then for any
    $n-\tau^{*}\geq\frac{2(c-p)+4\sqrt{p\log(3/\alpha)}-8\log(\alpha/3)}{\lVert\delta\rVert^{2}}=\Delta^{\prime}(\delta,c,\alpha,p)$
    we have
    $\mathbb{P}\Big{(}\lVert\frac{\sum_{t=\tau^{*}+1}^{n}y_{t}}{\sqrt{n-\tau^{*}}}\|^{2}<c\Big{)}\leq\alpha.$
  • Suppose the threshold $c_{n}$ is increasing, concave and its derivative $c^{\prime}_{n}$ well-defined for all integers $n$. Then for any $n-\tau\geq\Delta(\delta,c_{\tau^{*}},\alpha)+2\frac{c^{\prime}_{\tau^{*}}}{\lVert\delta\rVert^{2}}$ we have
    $\mathbb{P}\Big{(}\lVert\frac{\sum_{t=\tau^{*}+1}^{n}y_{t}}{\sqrt{n-\tau^{*}}}\|^{2}<c_{n}\Big{)}\leq\alpha.$
@@
We define $u_{n}=\sum_{t=\tau+1}^{n}\langle\frac{\delta}{\lVert\delta\rVert},\varepsilon_{t}\rangle$. Note that $u_{n}$ is $\mathcal{N}(0,1)$.
We get
$\left\lVert\frac{\sum_{t=\tau+1}^{n}y_{t}}{\sqrt{n-\tau}}\right\|^{2}$ $=$ $\left\lVert\sqrt{n-\tau}\delta+\frac{\sum_{t=\tau+1}^{n}\varepsilon_{t}}{\sqrt{n-\tau}}\right\rVert^{2}$
$=$ $(n-\tau)\lVert\delta\rVert^{2}+2\lVert\delta\lVert\sum_{t=\tau+1}^{n}\langle\frac{\delta}{\lVert\delta\rVert},\varepsilon_{t}\rangle+\left\lVert\frac{\sum_{t=\tau+1}^{n}\varepsilon_{t}}{\sqrt{n-\tau}}\right\rVert^{2}$
$\geq$ $(n-\tau)\lVert\delta\rVert^{2}+2\lVert\delta\lVert u_{n}$
$\geq$ $(\sqrt{n-\tau}\lVert\delta\rVert+u_{n})^{2}-u_{n}^{2}\,.$
Thus, we have:
$(n-\tau)\lVert\delta\rVert^{2}\geq\left(\sqrt{c_{n}+u_{n}^{2}}-u_{n}\right)^{2}$ $\iff$ $(\sqrt{n-\tau}\lVert\delta\rVert+u_{n})^{2}-u_{n}^{2}\geq c_{n}$ (26)
$\implies$ $\left\lVert\frac{\sum_{t=\tau+1}^{n}y_{t}}{\sqrt{n-\tau}}\right\|^{2}\geq c_{n}\,.$
Using the fact that for any $(a,b)$ in $\mathbb{R}^{2}$, $2a^{2}+2b^{2}\geq(a-b)^{2}$ we get
$(n-\tau)\lVert\delta\rVert^{2}\geq 2c_{n}+4u_{n}^{2}\implies(n-\tau)\lVert\delta\rVert^{2}\geq\left(\sqrt{c_{n}+u_{n}^{2}}-u_{n}\right)^{2}\,.$
In terms of probability events, we are looking for a delay $(n-\tau)$ such that the probability of detecting a change is greater than $1-\alpha$, this can be written as:
$1-\alpha\leq\mathbb{P}\Big{(}(n-\tau)\lVert\delta\rVert^{2}\geq 2c_{n}+4u_{n}^{2}\Big{)}\leq\mathbb{P}\Big{(}\left\lVert\frac{\sum_{t=\tau+1}^{n}y_{t}}{\sqrt{n-\tau}}\right\|^{2}\geq c_{n}\Big{)}\,.$
It is known that $u_{n}$ is $\mathcal{N}(0,1)$ and that $1-\alpha\leq\mathbb{P}(-2\log(\alpha/2)>u_{n}^{2})$ so we solve for a fixed threshold $c_{n}=c$: $-8\log(\alpha/2)\leq(n-\tau)\lVert\delta\rVert^{2}-2c$, leading to the relation:
$n-\tau\geq\frac{2c-8\log(\alpha/2)}{\lVert\delta\rVert^{2}}\,.$
For the second result, we proceed similarly but including a lower bound on $\lVert\frac{\sum_{i=\tau+1}^{n}\varepsilon_{i}}{\sqrt{n-\tau}}\rVert^{2}$ using Laurent and Massart [2000]
$\left\lVert\frac{\sum_{i=\tau+1}^{n}\varepsilon_{i}}{\sqrt{n-\tau}}\right\rVert^{2}\geq p-2\sqrt{p\log(3/\alpha)}\,.$
For the third result, we use the concavity property: $c_{n}\leq c_{n_{0}}+(n-n_{0})c^{\prime}_{n_{0}}$ in order to write:
$-8\log(\alpha/2)+2c_{n}\leq-8\log(\alpha/2)+2(c_{n_{0}}+(n-n_{0})c^{\prime}_{n_{0}})\leq(n-\tau)\lVert\delta\rVert^{2}\,.$
If the second inequality is true, the full line is true, which leads to the result with $n_{0}=\tau$. Notice that the bound $c_{n}$ is often chosen proportional to $\log(n)$ such that the delay found with constant $c_{n}$ is shifted by a constant when $\lVert\delta\rVert^{2}\geq\frac{1}{n_{0}}\frac{n-n_{0}}{n-\tau}$.
Appendix J Acknowledgements and Funding
This work was supported by EPSRC grants EP/Z531327/1, EP/W522612/1 and EP/N031938/1 and an ATIGE grant from Génopole. The IPS2 benefit from the support of Saclay Plant Sciences-SPS [ANR-17-EUR-0007]